stan has 52 quarters and nickels combined. if he has three times as many nickels as quarters, how much money does he have?
Found 2 solutions by stanbon, checkley75:
Answer by stanbon(75887)

stan has 52 quarters and nickels combined. if he has three times as many nickels as quarters, how much money does he have?
------------
Let # of quarters be "x"
------------
# of nickels is "3x"
-------------
EQUATIOn:
x + 3x = 52
4x = 52
x = 13 (# of quarters; value of these is 25*13 = $3.25
3x = 39 (# of nickels; value of these is 39*5 = $1.95
---------------
He has $5.20
=================
Cheers,
Stan H.

------------

Answer by checkley75(3666)

3N+N=52
4N=52
N=52/4
N=13 OR 3*13=39 NICKLES & 13 QUARTERS.
39*5+13*25=195+325=$5.20 TOTAL AMOUNT OF MONEY.